OpenStudy (anonymous):

Select all that justify the following statement! 8+1/8+(-1)+5=5 1. commutative - addition 2. inverse - addition 3. associative - multiplication 4. symmetric 5. commutative - multiplication 6. associative - addition 7. inverse - multiplication

OpenStudy (anonymous):

can you use some parenthesis? is that ((8+1)/8) -1 + 5 = 5? or 8 + (1/8) -1 + 5 = 5 you gotta use the parens so i know the order.

OpenStudy (anonymous):

(1/8) lol sorry

OpenStudy (anonymous):

8+(1/8)+(-1)+5=5 8 + 0.125 - 1 + 5 = 5 12.125 = 5 I don't think any statements can justify that. Is there a "none" answer?